# Iband Measurements in a Sample of Hosts of 9 SN Ia Match HST Cepheids

### v1

@@ 6 Discussion
We measure F090W TRGB-based distances to NGC 1448, NGC 1559, NGC 2525, NGC 3370, NGC 3447, NGC 5584, and NGC 5643 using JWST NIRCam observations and test their consistency with HST Cepheid-based distances to those same galaxies. We find excellent agreement between the two independent sets of distance measurements, which were measured between the same hosts and anchor but using data taken with different telescopes and distance indicators.
@@
It will be important to also scrutinize the difference in direct measurements that arise from different treatments of SNe Ia. Scolnic et al. (2023) find that the addition of peculiar flow corrections and cross-calibration of datasets which are included in the Pantheon and Pantheon+ analyses increase measurements each by 0.5 and 1.1 km/s/Mpc from the Freedman (2021a) analysis, bringing it closer in line with the value obtained from Cepheids. An under recognized source of variations in determinations of is also the make-up of the SN Ia calibration sample which can produce fluctuations of 1-2 km/s/Mpc due to small sample statistics.
@@
Crosschecks between standard candles can be made more comprehensive with additional, independent standard candles such as Miras and carbon stars. The carbon star method, also called the J-region Asymptotic Giant Branch (JAGB), originated in the 1980s (Richer, 1981; Richer et al., 1984, 1985; Pritchet et al., 1987; Cook et al., 1986) and has since been revived and further pioneered in the 2000s (Battinelli & Demers, 2005; Ripoche et al., 2020; Madore & Freedman, 2020; Freedman & Madore, 2020; Parada et al., 2021, 2023; Zgirski et al., 2021; Lee et al., 2021b, a; Madore et al., 2022; Lee et al., 2022; Lee, 2023; Lee et al., 2024; Li et al., 2024). Although the JAGB method still requires further development and standardization (see discussions of asymmetric luminosity functions and metallicity effects in Parada et al. 2021, 2023; Li et al. 2024), the great luminosities of carbon stars in the near-infrared allow the JAGB to reach farther than the I and J-band TRGB, potentially allowing for a more comprehensive crosscheck of Cepheid variables via galaxy distances.
@@
The TRGB measurements made here can benefit from further analysis, such as a future contrast ratio calibration (see CATs, Wu et al. (2023); Scolnic et al. (2023); Li et al. (2023b) or via modelling of the luminosity function (Makarov et al., 2006; Anand et al., 2019b, 2024a) instead of edge-detection algorithms (Anand et al., in prep). However, even without these additions, the consistency of the TRGB-based distance measurements with those of Cepheid-based distances measurement do not show evidence of a bias or systematic offset in Cepheid distances that would resolve the Hubble Tension. Similar analyses in the future will also benefit from further studies investigating the effects of red giant diversity to improve the accuracy of TRGB distances (such as in Anderson et al., 2024; Koblischke & Anderson, 2024). Future observations of TRGB fields with HST (for instance, from GO-17520; Breuval et al. 2023), JWST, and the Roman Space Telescope will provide more opportunities to scrutinize the second rung of the distance ladder by providing a more extensive set of observations that can be used to compile a more comprehensive comparison between these distance indicators. Specifically, wide-field JWST observations within the anchor galaxy NGC 4258 will help reduce systematic uncertainties with regards to the absolute calibration of the TRGB, and deep JWST observations of galaxies with the highest levels of mismatch between HST Cepheid and HST TRGB distances (e.g. NGC 3021, NGC 4038/9; Jang & Lee 2017b) will help elucidate any individual discrepancies between the two techniques. Lastly, a fully independent distance ladder from the traditional Cepheid+SNe Ia route is also under construction, via the usage of the TRGB combined with Surface Brightness Fluctuations as the final rung (Anand et al., 2024b).

### v2

@@ 6 Discussion
We measure F090W TRGB-based distances to NGC 1448, NGC 1559, NGC 2525, NGC 3370, NGC 3447, NGC 5584, NGC 5643, and NGC 5861 using JWST NIRCam observations and test their consistency with HST Cepheid-based distances to those same galaxies. We find excellent agreement between the two independent sets of distance measurements, which were measured between the same hosts and anchor but using data taken with different telescopes and distance indicators.
@@
It will be important to also scrutinize the difference in direct measurements that arise from different treatments of SNe Ia. Scolnic et al. (2023) find that the addition of peculiar flow corrections and cross-calibration of datasets which are included in the Pantheon and Pantheon+ analyses increase measurements each by 0.5 and 1.1 km/s/Mpc from the Freedman (2021a) analysis, bringing it closer in line with the value obtained from Cepheids. An under recognized source of variations in determinations of is also the make-up of the SN Ia calibration sample which can produce fluctuations of 12 km/s/Mpc due to small sample statistics.
@@
Crosschecks between standard candles can be made more comprehensive with additional, independent standard candles such as Miras and carbon stars. The carbon star method, also called the J-region Asymptotic Giant Branch (JAGB), originated in the 1980s (Richer, 1981; Richer et al., 1984, 1985; Pritchet et al., 1987; Cook et al., 1986) and has since been revived and further pioneered in the 2000s (Battinelli & Demers, 2005; Ripoche et al., 2020; Madore & Freedman, 2020; Freedman & Madore, 2020; Parada et al., 2021, 2023; Zgirski et al., 2021; Lee et al., 2021b, a; Madore et al., 2022; Lee et al., 2022; Lee, 2023; Lee et al., 2024; Li et al., 2024). Although the JAGB method still requires further development and standardization (see discussions of asymmetric luminosity functions and metallicity effects in Parada et al. 2021, 2023; Li et al. 2024), the great luminosities of carbon stars in the near-infrared allow the JAGB to reach farther than the I- and J-band TRGB, potentially allowing for a more comprehensive crosscheck of Cepheid variables via galaxy distances.
@@
The TRGB measurements made here can benefit from further analysis, such as a future contrast ratio calibration (see CATs; Wu et al. 2023; Scolnic et al. 2023; Li et al. 2023b) or via modelling of the luminosity function (Makarov et al., 2006; Anand et al., 2019b, 2024a) instead of edge-detection algorithms (Anand et al., in prep). However, even without these additions, the consistency of the TRGB-based distance measurements with those of Cepheid-based distances measurement do not show evidence of a bias or systematic offset in Cepheid distances that would resolve the Hubble Tension. Similar analyses in the future will also benefit from further studies investigating the effects of red giant diversity to improve the accuracy of TRGB distances (such as in Anderson et al., 2024; Koblischke & Anderson, 2024). Future observations of TRGB fields with HST (for instance, from GO-17520; Breuval et al. 2023), JWST, and the Roman Space Telescope will provide more opportunities to scrutinize the second rung of the distance ladder by providing a more extensive set of observations that can be used to compile a more comprehensive comparison between these distance indicators. Specifically, wide-field JWST observations within the anchor galaxy NGC 4258 will help reduce systematic uncertainties with regards to the absolute calibration of the TRGB, and deep JWST observations of galaxies with the highest levels of mismatch between HST Cepheid and HST TRGB distances (e.g. NGC 3021, NGC 4038/9; Jang & Lee 2017b) will help elucidate any individual discrepancies between the two techniques. Lastly, a fully independent distance ladder from the traditional Cepheid+SNe Ia route is also under construction, via the usage of the TRGB combined with Surface Brightness Fluctuations as the final rung (Anand et al., 2024b).